# SHuBERT: Self-Supervised Sign Language Representation Learning via Multi-Stream Cluster Prediction

### v1

@@ Abstract
Sign language processing has traditionally relied on task-specific models, limiting the potential for transfer learning across tasks. We introduce SHuBERT (Sign Hidden-Unit BERT), a self-supervised transformer encoder that learns strong representations from approximately 1,000 hours of American Sign Language (ASL) video content. Inspired by the success of the HuBERT speech representation model, SHuBERT adapts masked prediction for multi-stream visual sign language input, learning to predict multiple targets for corresponding to clustered hand, face, and body pose streams. SHuBERT achieves state-of-the-art performance across multiple benchmarks. On sign language translation, it outperforms prior methods trained on publicly available data on the How2Sign (+0.7 BLEU), OpenASL (+10.0 BLEU), and FLEURS-ASL (+0.3 BLEU) benchmarks. Similarly for isolated sign language recognition, SHuBERT’s accuracy surpasses that of specialized models on ASL-Citizen (+5%) and SEM-LEX (+20.6%), while coming close to them on WLASL2000 (-3%). Ablation studies confirm the contribution of each component of the approach.
@@ 1 Introduction
Sign languages are used by millions of individuals worldwide, but the development of automatic sign language processing systems lags far behind that of systems for spoken and written language [62]. Unlike speech or text processing, sign language presents unique challenges due to its multi-channel nature, combining manual gestures, facial expressions, and body movements, which can be quick and highly coarticulated [3].
Existing approaches to sign language processing have typically relied on models designed and trained for specific tasks, such as translation from signed to written languages [6, 33, 63], isolated sign recognition [27, 25], and fingerspelling detection and recognition [49, 14, 17]. These approaches often require large amounts of annotated data, which is particularly scarce for sign languages. This limitation has historically constrained the performance and scalability of sign language processing systems.
@@
Figure 1: Comparison between our results using fine-tuned SHuBERT and results of the previous state-of-the-art task-specific models on a suite of multiple tasks, datasets, and metrics. Note: The orange shade does not represent a single model but a collection of the previous SOTA results on these tasks, datasets, and metrics. The SHuBERT-based results improve on all but one task-specific SOTA model in our experiments. See Sec. 4 for more experimental details.
@@
Figure 2: SHuBERT pre-training. (a) We locate a set of landmarks in each frame of the input video using MediaPipe [35], with inter-frame interpolation to fill in missing landmarks. From these, we extract the upper body pose, crop the hand and face regions, and blur and partially mask the face crop is blurred for a measure of privacy. (b) We extract body pose and DINOv2 features for the hands and face, yielding a four-stream representation (two hands, face, body pose) for each frame. (c) We assign the feature vectors for frame $t$ to cluster indices using pre-computed $k$-means clusters, yielding assignments $(f_{t},l_{t},r_{t},b_{t})\in[k]^{4}$ for face, left and right hand and body pose, respectively. (d) We partially mask the features, which form the input to a transformer encoder. The output of the transformer is fed to a linear classifier (e) predicting the cluster assignments for each frame, $(\widehat{f}_{t},\widehat{l}_{t},\widehat{r}_{t},\widehat{b}_{t})$. We train the linear layer and the encoder using the log-loss of the predicted vs. true assignments for the masked tokens.
Recent advances in self-supervised learning of representations for natural language and speech have produced representations that perform well across a wide range of tasks. For example, encoder models like BERT for written language [12, 31] and HuBERT for speech [20] are learned from unlabeled data by predicting masked portions of the input. Self-supervised learning seems particularly relevant for sign languages, for which annotated data are particularly scarce. But the unique visual properties of sign languages, combining interacting fine-grained hand and facial gestures and global body motion, demand a specialized approach.
@@
In this work, we present SHuBERT (Sign Hidden-Unit BERT) (see  Fig. 2), a self-supervised representation learning approach that adapts the masked prediction paradigm to the characteristics of sign language video. SHuBERT learns contextual, transformer-based representations of video frames by predicting cluster assignments of multiple masked video streams of the hands, face, and body pose features. We demonstrate that the resulting learned representations transfer effectively to multiple downstream tasks, achieving state-of-the-art performance on several sign language translation (SLT) benchmarks, as well as on multiple isolated sign recognition (ISLR) benchmarks. Our results suggest that a self-supervised learning approach to sign language representation learning is not only possible, but can outperform specialized architectures while requiring less task-specific annotation. We conclude that SHuBERT is a promising candidate for the role of foundation model [4] for sign language video processing.
@@ 2 Related Work
Advances in sign language processing have often been based on inspiration from methods found to be successful for other video domains and for text. This section first examines influential pre-training methods in text and speech processing that inform SHuBERT, and then explores specific developments in sign language understanding, including both supervised and self-supervised approaches.
@@ 2.1 Pre-Training for Text And Speech
Pre-training has become a cornerstone of modern language processing across modalities. In natural language (text) processing, encoder models like BERT [12] and its variants (e.g., RoBERTa [34] and ALBERT [31]), based on masked language modeling, have served as dependable representations that can be used and adapted for a variety of language understanding tasks [4]. Pre-trained encoder-decoder models, like T5 [43], unify multiple text-to-text tasks in a single model. Finally, decoder-only models like the GPT series [42] have demonstrated the power of auto-regressive pre-training, when very large-scale training data is available and the downstream tasks can be framed in terms of auto-regressive generation.
@@
In speech processing, self-supervised learning approaches have taken inspiration from text encoder models while addressing the unique challenges posed by continuous audio signals, which have no inherent segmentation into tokens nor a pre-defined token vocabulary. For example, Hidden-Unit BERT (HuBERT) [20] adapted the objective of BERT by adding an offline clustering step to provide pseudo-labels for masked prediction. Other approaches include wav2vec 2.0 [2], which uses a contrastive objective, and AV-HuBERT [51] which extends HuBERT to audiovisual speech (i.e., speech along with video of the speaker’s mouth). Signed languages share similar challenges to speech, with no pre-existing token lexicon and variable-length units (gestures) with no explicit boundaries, and our approach takes inspiration from HuBERT.
2.2 Pre-Training for Sign Language Understanding
Sign language processing research has explored both supervised and self-supervised pre-training approaches for individual tasks.
@@ 2.2 Pre-Training for Sign Language Understanding
For translation of sign language video, the supervised pre-training approach has focused on first (pre-)training a translation model on a large (but often noisy) out-of-domain dataset, followed by fine-tuning on a smaller in-domain dataset. This type of pre-training leverages large collections of annotated data, with some systems [55, 54, 52, 63] trained on up to 6,600 hours of sign language content [55] to achieve state-of-the-art performance. However, much of this data remains private, including the largest dataset used by [63]. In the self-supervised line of work, SSVP-SLT [44] adapts masked autoencoders for sign language video, achieving competitive performance on the sign language translation task. These approaches, while effective, often involve substantial computational resources: the supervised model of [63] was trained on 128 TPU-v3 chips for 20 days, while the MAE-like self-supervised approach of [44] used 64 A100 GPUs for 14 days.
Other lines of work (e.g., [8]) have use a pre-trained S3D model [60], which requires the video sequence to be segmented into chunks, with each chunk treated as independent by the model. This form of pre-training therefore doesn’t capture long temporal context. SHuBERT instead learns contextual frame representations and can therefore operate on long video sequences directly.
For isolated sign language recognition, several self-supervised approaches have shown promise. SignBERT [22] is a self-supervised model that treats hand poses as visual tokens and learns via masked prediction of hand joints. An extension of this approach is SignBERT+ [23], which is a self-supervised pre-training approach for sign language that focuses on hand pose modeling with a specialized hand-model-aware decoder. In contrast, SHuBERT models all components of sign language by treating it as a multi-stream signal, simultaneously modeling the face, hands, and body pose. Another recent study compared multiple self-supervised techniques for ISLR, finding masked autoencoders (MAE) particularly effective [45].
@@ 2.3 Multi-Stream Models of Sign Language
Another recent translation approach, SignMusketeers [19], found that it is possible to achieve similar performance to SSVP-SLT while using less than 5% of the compute, by using a multi-stream model of the hands, face, and body. SHuBERT also adopts the idea of multiple streams. However, unlike prior multi-stream approaches, SHuBERT learns a self-supervised representation that learns from the multiple streams jointly and that performs well on multiple tasks. Several previous methods have taken advantage of the observation that sign language naturally decomposes into multiple streams of hand, face, and body motions. In addition to the aforementioned SignMusketeers, prior work has developed multi-stream models for fingerspelling recognition (which combines hand and mouthing gestures) [48], sign language translation  [50, 7, 65, 9], and isolated sign language recognition [41, 24]. SHuBERT also uses a decomposition into multiple streams of information, but learns a representation in a self-supervised way that combines information from all of them.
@@ 3 Sign Hidden-Unit BERT (SHuBERT)
SHuBERT is a transformer encoder [56] that learns contextualized representations of sign language video frames through self-supervised learning. The pre-training approach is outlined in Fig. 2. In the following sections we describe the video features used in SHuBERT (Sec. 3.1) and the self-supervised training approach (Sec. 3.2).
@@ 3.1 Multi-Stream Feature Pre-Processing
We use the MediaPipe Hand Landmarker [^1] https://ai.google.dev/edge/mediapipe/solutions/vision/hand_landmarker model, which has hand detection accuracy 95%. Upon inspection, we find that the majority of the remaining 5% of “failed” detections occur when the hands are outside the frame. For these cases, we interpolate from the nearest frames with successful detections. Dilated bounding boxes for the detected hand landmarks (for both left and right hand) are cropped and resized to 224×224.
@@
The signer’s face contains important non-manual markers for sign languages [5]. Previous approaches either use the full face, compromising privacy [19], or completely blur the whole face in an attempt to protect privacy, potentially losing essential non-manual markers [44]. Our design attempts to balance the need to preserve linguistic information conveyed by the face with the goal of enhancing the signer’s privacy. We identify the whole face, mouth and eye regions in the frame from the relevant MediaPipe facial landmarks. The face pixels are greyed out except for the pixels in the mouth and eye regions. We then apply Gaussian blur to the entire face region, resize it to 224×224, and feed it into our DINO-based extractor.
Self-Supervised Image Feature Extractor
For face and hand image crops, while most works use pose, we use DINOv2 [38] for feature extraction, which has proven to be a robust feature extractor for sign language in previous work [59, 19]. An additional benefit of DINOv2 representations is that they yield meaningful clusters when quantized [64], an important property since SHuBERT training targets are clustered input features. While most other works use keypoints estimation tools, it has shown to have some weakness to adequately capture sign language handshapes [36] and facial expressions [30].
To adapt the general image feature extractor to a face feature extractor for sign language, we use 5 million random face crops from videos in the YouTube-ASL [55] dataset to fine-tune DINOv2 for 1 epoch. We do the same for the hand feature extractor, using 5 million random hands (mixing left and right hand crops) from YouTube-ASL [55]. For the face and hand streams, crop regions of interest (ROIs) are processed through the face or hand fine-tuned DINOv2 models, yielding a 384-dimensional feature vector per crop.
Body Posture
For body pose, we extract seven key upper body landmarks (nose, shoulders, elbows, and wrists) and normalize their coordinates relative to the signing space, resulting in a compact 14-dimensional pose vector.
@@ 3.2 Self-Supervised Training of SHuBERT
Figure 4: Three masking strategies for sequence masking. Empirically, random masking results in the best translation performance (see Tab. 8).
Multi-Stream Input
We combine the four feature vectors described above—a 384-dimensional feature vector for each hand, a 384-dimensional face feature vector, and a 14-dimensional pose feature vector—resulting in 1,166 dimensions. We extract these features at every second frame.
Specifically, each channel first undergoes independent layer normalization to handle the varying scales of the different streams. Each channel’s normalized features are then projected through a separate linear layer to 256-dimension each. These projected channel-specific representations are concatenated to create a unified frame-level representation, which captures information from all modalities. This combined representation is then input to the transformer encoder to model temporal relationships across frames.
@@
Inspired by prior work in self-supervised learning in speech [2, 20], SHuBERT learns by predicting masked elements of the input feature sequence, given the observed (unmasked) data. We use a masking strategy designed for multi-channel sign language input. We consider three types of masking, illustrated in Fig. 4: channel masking, which masks entire channels (e.g., all face and left hand features in a video) to learn cross-channel dependencies; time masking, which masks all channels at selected temporal positions (e.g., face, hand, and body pose in frames 20-40 in a given video); and random masking, which independently masks random frames in each channel. Based on our experiments comparing these strategies (Sec. 4.5), we ultimately chose random masking as our final approach. After replacing the masked inputs with a mask token, all of the feature frames are fed through the transformer encoder to produce contextualized representations.
@@
For the masked input elements, the representation at the last transformer layer is used to predict target pseudo-labels. Inspired by HuBERT [20], we use offline $k$-means clustering (separately for each of the four channels) to create discrete target pseudo-labels. Target labels are extracted independently for each channel, resulting in four separate cluster assignments per frame. A separate prediction head is used for each channel to predict the corresponding cluster assignment for masked regions.
The training objective is a cross-entropy loss between the target and predicted cluster assignments, computed over the masked positions.

### v3

@@ Abstract
Sign language processing has traditionally relied on task-specific models, limiting the potential for transfer learning across tasks. Pre-training methods for sign language have typically focused on either supervised pre-training, which cannot take advantage of unlabeled data, or context-independent (frame or video segment) representations, which ignore the effects of relationships across time in sign language. We introduce SHuBERT (Sign Hidden-Unit BERT), a self-supervised contextual representation model learned from approximately 1,000 hours of American Sign Language video. SHuBERT adapts masked token prediction objectives to multi-stream visual sign language input, learning to predict multiple targets corresponding to clustered hand, face, and body pose streams. SHuBERT achieves state-of-the-art performance across multiple tasks including sign language translation, isolated sign language recognition, and fingerspelling detection.
SHuBERT: Self-Supervised Sign Language Representation Learning
via Multi-Stream Cluster Prediction
Shester Gueuwou1, Xiaodan Du1, Greg Shakhnarovich1, Karen Livescu1, Alexander H. Liu2 1TTI-Chicago, 2MIT CSAIL {shesterg,xdu,greg,klivescu}@ttic.edu, alexhliu@mit.edu http://shubert.pals.ttic.edu
@@ 1 Introduction
Sign language presents unique challenges compared to other language modalities, because of the relative scarcity of data and its multi-channel nature, combining manual, facial, and other body movements, which can be quick and highly coarticulated Bellugi and Fischer (1972). Existing approaches to sign language processing have typically relied on models designed and trained for specific tasks, such as sign language translation (SLT) from signed to written languages Camgoz et al. (2018); Shi et al. (2022); Zhang et al. (2024), isolated sign language recognition (ISLR) Kezar et al. (2023), and fingerspelling detection and recognition Shi et al. (2019); Fayyazsanavi et al. (2024); Georg et al. (2024). Pre-training approaches allow for pooling data across tasks, and several pre-training methods have been successful for sign language tasks Uthus et al. (2023); Rust et al. (2024). However, these have typically focused on either supervised pre-training, which cannot take advantage of unlabeled data, context-independent (frame or video segment) representations, which ignore the effects of relationships across time in sign language, or contextual representations of only some aspects of sign language (see Sec. 2). These limitations have historically constrained the performance and scalability of sign language processing systems.
@@
Figure 1: Comparison between our results using fine-tuned SHuBERT and results of the previous state-of-the-art task-specific models on a suite of tasks, datasets, and metrics. Note: The orange shade does not represent a single model but a collection of the previous SOTA results for models trained on public data. The SHuBERT-based results improve on all but one task-specific SOTA model. See Sec. 4 for details.
@@
Figure 2: SHuBERT pre-training. (a) We locate a set of landmarks in each frame of the input video using MediaPipe Lugaresi et al. (2019), with inter-frame interpolation to fill in missing landmarks. From these, we extract the upper body pose, crop the hand and face regions, and blur and partially mask the face crop for a measure of privacy and robustness. (b) We use DINOv2 Oquab et al. (2023) to extract features for the hands and face, yielding a four-stream representation (two hands, face, body pose) for each frame. (c) We assign the feature vectors for frame $t$ to cluster indices using pre-computed $k$-means clusters, yielding assignments $(f_{t},l_{t},r_{t},b_{t})\in[k]^{4}$ for face, left and right hand, and body pose, respectively. (d) We partially mask the features, and the masked features form the input to the transformer encoder. (e) We train SHuBERT to predict the cluster assignments for each masked input frame, $(\widehat{f}_{t},\widehat{l}_{t},\widehat{r}_{t},\widehat{b}_{t})$.
The success of self-supervised representations for written and spoken language, such as BERT for written language Devlin et al. (2019) and HuBERT for speech Hsu et al. (2021), has yet to be realized for sign language. Self-supervised learning seems particularly relevant for sign language, for which annotated datasets are scarce. But the unique multi-channel and other visual properties of sign languages suggest a specialized approach.
@@
In this work, we present SHuBERT (Sign Hidden-Unit BERT) (Fig. 2), a self-supervised representation learning approach that learns contextual frame representations for all sign language channels jointly. SHuBERT adapts the masked prediction paradigm of BERT and HuBERT to the characteristics of sign language video, and learns by predicting cluster assignments of multiple masked feature streams representing the hands, face, and body pose. The learned representations transfer effectively to multiple sign language understanding tasks, achieving state-of-the-art performance on several SLT benchmarks, multiple ISLR benchmarks, and fingerspelling detection, and improving over specialized models for each task (Fig. 1).
@@ 2 Related Work
Sign language understanding (recognition and translation) tasks have received increasing attention in the last few years Camgoz et al. (2018); Shi et al. (2022); Lin et al. (2023); Kezar et al. (2023). For translation, early work mainly focused on gloss-based methods, which rely on (the rare and small) datasets with manually labeled glosses Camgoz et al. (2018). More recent work has turned to larger and more naturalistic datasets without gloss labels. The most commonly used datasets are in American Sign Language (ASL) Duarte et al. (2021); Shi et al. (2022); Uthus et al. (2023), German Sign Language (DGS) Camgoz et al. (2018), British Sign Language Albanie et al. (2021), and Chinese Sign Language Zhou et al. (2021a). Of these, recent ASL datasets are the most naturalistic, and include large quantities of natively produced sign language (rather than translated from a spoken language, which has properties of “translationese” Desai et al. (2024b)). For this reason we focus on ASL data and tasks, but our approach is applicable and extensible to any sign language.
@@ 2.1 Pre-Training for Text And Speech
Pre-training is a cornerstone of modern language processing across modalities. For written language, encoder models like BERT Devlin et al. (2019) and its variants (e.g., Liu (2019); Lan et al. (2020)), based on masked language modeling, have served as dependable representations for language understanding tasks. In speech processing, self-supervised learning approaches Mohamed et al. (2022) have taken inspiration from text encoder models while addressing the unique challenges posed by continuous audio, which has no inherent segmentation into tokens nor a pre-defined token vocabulary. For example, Hidden-Unit BERT (HuBERT) Hsu et al. (2021) adapts BERT by adding an offline clustering step to provide pseudo-labels for masked prediction. Such self-supervised representations, combined with task-specific fine-tuning, remain the state of the art for many speech tasks.
Sign languages share similar challenges to speech, with no pre-existing token lexicon and variable-length units (gestures) with no explicit boundaries, and our approach takes inspiration from HuBERT. However, sign language video has its own unique challenges: the many sources of variation (signer appearance, background, lighting, camera angles), the multiple streams of gestures (hands, face, body), the high dimensionality of video, and the relative dearth of data. These challenges are addressed in SHuBERT by focusing on the relevant streams (via pose tracking) combined with multi-stream masking and clustering.
2.2 Pre-Training for Sign Language
Supervised pre-training.
@@ 2.2 Pre-Training for Sign Language
For translation of sign language video, the supervised pre-training approach has focused on (pre-)training a translation model on a large (but often noisy) out-of-domain dataset, followed by fine-tuning on a smaller in-domain dataset. This type of pre-training leverages large collections of annotated data, with some systems Uthus et al. (2023); Tanzer and Zhang (2024); Tanzer (2024a); Zhang et al. (2024) trained on up to 6,600 hours of sign language content Uthus et al. (2023) to achieve state-of-the-art performance. However, much of this data remains private. In addition, these approaches often involve substantial computational resources: The supervised model of  Zhang et al. (2024), for example, was trained on 128 TPU-v3 chips for 20 days. Jiao et al. (2024) propose an alternative pre-training approach that greatly improves efficiency by using pose information only (rather than image pixels); however, this approach pre-trains and fine-tunes on the same training data. Unlike these approaches, Uni-Sign Li et al. (2025) is a supervised pre-training approach, based on mT5 Xue et al. (2021), that has been applied to multiple tasks including both translation and ISLR. Like all supervised methods, these approaches can not take advantage of available unlabelled sign language data.
Self-supervised pre-training.
Previous work has compared multiple context-independent self-supervised techniques for ISLR, finding masked autoencoders (MAE) particularly effective Sandoval-Castaneda et al. (2023). SSVP-SLT Rust et al. (2024) adapts MAEs for large-scale sign language pre-training, achieving competitive performance on ASL-to-English translation. This approach is computationally demanding, using 64 A100 GPUs for 14 days, and takes a maximum of 128 input frames ($\sim$8 seconds) at a time so is unable to model longer-term dependencies. Other lines of work on SLT (e.g., Chen et al. (2022a)) have used a pre-trained S3D model Xie et al. (2018), which requires the video sequence to be segmented into chunks, with each chunk treated as independent. All of these approaches learn context-independent representations of individual frames or video segments, whereas SHuBERT learns contextual frame representations and can operate on long video directly.
The only previous self-supervised approach of which we are aware for contextual sign representation learning is SignBERT+ Hu et al. (2023), which extends the earlier SignBERT Hu et al. (2021). This approach learns a representation specifically for hand poses, via masked reconstruction of hand joints, and has strong results on ISLR, continuous (gloss-based) sign recognition, and sign translation on the RWTH-PhoenixT German Sign Language dataset Camgoz et al. (2018). However, this approach is inherently limited by not modeling the face and global body pose, and the results are obtained by combining SignBERT+ with a dataset-specific image pixel (RGB) representation. In addition, SignBERT+ is pre-trained on the union of datasets on which it is tested; that is, it is exposed to the fine-tuning data during pre-training. In contrast, SHuBERT models all components of sign language jointly, and is pre-trained on data that is disjoint from the fine-tuning data for the downstream tasks.
@@ 2.3 Multi-Stream Models of Sign Language
Several previous methods have taken advantage of the observation that sign language naturally decomposes into multiple streams of hand, face, and body motions. [^1] The term “multi-stream” has been used in different senses in prior work. For example, DSTA-SLR Hu et al. (2024) creates multiple streams consisting of different geometric representations of the same skeleton data, while we are concerned with streams that correspond to different body parts. For example, prior work incudes multi-stream models for fingerspelling recognition (which combines hand and mouthing gestures) Shi (2023), SLT Camgoz et al. (2020); Zhou et al. (2021b); Chen et al. (2022b); Shi et al. (2022); Gueuwou et al. (2025), and ISLR Pu et al. (2016); Jiang et al. (2021). This factorization into multiple streams can enable dramatic improvements in data and compute efficiency over single-stream models that use the full image Gueuwou et al. (2025).
Like this prior work, SHuBERT also adopts the idea of multiple streams. However, unlike prior approaches, SHuBERT learns a self-supervised representation from the multiple streams jointly that performs well on multiple tasks.
@@ 3 Sign Hidden-Unit BERT (SHuBERT)
SHuBERT is a transformer encoder Vaswani et al. (2017) that learns contextualized representations of sign language video frames through self-supervised learning. The pre-training approach is outlined in Fig. 2. In the following sections we describe the video features used in SHuBERT (Sec. 3.1) and the self-supervised training approach (Sec. 3.2).
@@ 3.1 Multi-Stream Feature Pre-Processing
We use the MediaPipe Hand Landmarker [^2] https://ai.google.dev/edge/mediapipe/solutions/vision/hand_landmarker model, which has hand detection accuracy $\sim$95on OpenASL. [^3] Estimated from a small-scale experiment on 100 OpenASL videos, with manual verification of detections. Upon inspection, we find that the majority of the remaining 5% of “failed" detections occur when the hands are outside the frame. For these cases, we interpolate from the nearest frames with successful detections. Dilated bounding boxes for the detected hand landmarks (for both left and right hand) are cropped and resized to 224×224.
@@
The signer’s face contains important non-manual markers for sign languages Bragg et al. (2019). Previous approaches either use the full face, compromising privacy Gueuwou et al. (2025), or blur the whole face in an attempt to protect privacy, potentially losing essential non-manual markers Rust et al. (2024). Our design attempts to balance the need to preserve linguistic information with the goal of enhancing privacy. We identify the whole face, mouth and eye regions in the frame from the relevant MediaPipe facial landmarks. The face pixels are greyed out except for the pixels in the mouth and eye regions. We then apply Gaussian blur to the entire face region and resize it to 224×224.
Image Feature Extractor for Hands and Face
We use DINOv2 Oquab et al. (2023), which has proven successful in previous sign language work Wong et al. (2024); Gueuwou et al. (2025), as the feature extractor for face and hand image crops. An additional benefit of DINOv2 representations is that they yield meaningful clusters after quantization Zheng et al. (2024), which is an important property since SHuBERT training targets are clustered input features. While many other prior approaches use keypoint estimation tools, these have some weaknesses in capturing handshapes Moryossef et al. (2021) and facial expressions Kuznetsova and Kimmelman (2024).
To adapt the general image feature extractor to a face feature extractor for sign language, we randomly sample 5 million face crops from videos in the YouTube-ASL Uthus et al. (2023) dataset and use them for continued pre-training of DINOv2 for 1 epoch. We do the same for the hand feature extractor, using 5 million randomly sampled hands (mix of left and right hand crops) from YouTube-ASL. For both the face and hand streams, crop regions of interest (ROIs) are processed through the face or hand fine-tuned DINOv2 models, yielding a 384-dimensional feature vector per crop, which we denote $\mathbf{x}^{f}_{t},\mathbf{x}^{l}_{t},\mathbf{x}^{r}_{t}\in\mathbb{R}^{384}$ for the face, left hand, and right hand features respectively.
Body Pose
For (coarse) body pose, we extract seven key upper body landmarks (nose, shoulders, elbows, and wrists) and normalize their coordinates relative to the signing space, resulting in a compact 14-dimensional pose vector, $\mathbf{x}^{b}_{t}\in\mathbb{R}^{14}$.
@@ 3.2 Self-Supervised Training of SHuBERT
Figure 4: Three strategies for sequence masking. Each color-coded row corresponds to one of the four channels (face, right hand, left hand, body).
For each of the four channels (left/right hand, face, and pose), we layer-normalize the extracted features (producing zero mean and unit variance across all dimensions for each feature vector) and linearly project them to 256 dimensions, producing a 1024-dimensional input feature vector per frame. This joint representation of the four streams is masked (see below) and input to the transformer encoder. The output of the transformer is one representation vector per input frame, $y_{t}$, which may take into account information from the entire length of the input video, and from all four input channels.
@@
SHuBERT learns by predicting masked elements of the input feature sequence, given the observed (unmasked) data. We use a masking strategy designed for multi-channel sign language input. We consider three types of masking, illustrated in Fig. 4: channel masking, which masks entire channels (e.g., all face and left hand features in a video) to learn cross-channel dependencies; time masking, which masks all channels at selected temporal positions (e.g., face, hand, and body pose in frames 20-40 in a given video); and random masking, which independently masks random small frame spans in each channel. Based on our experiments comparing these strategies (Appendix A), we ultimately chose random masking.
@@
We use offline $k$-means clustering (separately for each of the four channels) to create discrete target pseudo-labels $(f_{t},l_{t},r_{t},b_{t})$ for the face, left hand, right hand, and body pose respectively. The transformer output vector for each frame is fed to four linear classifiers (one per channel) to predict the cluster assignments for masked channels for that frame.
As an example, suppose that face and body pose channel features for frame $t$, respectively $\mathbf{x}^{f}_{t}\in\mathbb{R}^{384}$ and $\mathbf{x}^{b}_{t}\in\mathbb{R}^{14}$, are masked. The $k$-means cluster assignments for these masked feature vectors are, respectively, $f_{t}$ and $b_{t}$, each a number between 1 and $k$. The classifier predicts, from the output vector $y_{t}$ for frame $t$, labels $\widehat{f}_{t}$ and $\widehat{b}_{t}$. The training objective is a cross-entropy loss between the target and predicted cluster assignments for the masked positions. The unmaksed positions are not included in the loss (but of course influence the predictions for the masked ones).